OpenStudy (anonymous):

solve the equation 10^x-3=100^4x-5

OpenStudy (amistre64):

divide by 10 wait, those are exponents ....

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

10^(x-3) ?? or (10^x) -3 ?

OpenStudy (anonymous):

the first

OpenStudy (amistre64):

log both sides: log10(10^(x-3)) = log10(100^(4x-5)) (x-3) = log10(100) (4x-5) x-3 = 2(4x-5) x - 3 = 8x - 10 10-3 = 8x-x 7 = 7x....x = 1

OpenStudy (amistre64):

10^2 = 100^1

OpenStudy (anonymous):

is that the answer?

OpenStudy (amistre64):

the answer is: x = 1

OpenStudy (anonymous):

oh yeah that makes sense thankyou

OpenStudy (amistre64):

youre welcome :)